OpenStudy (anonymous):

please prove this identity. tan +sec=1/sec*tan..

OpenStudy (anonymous):

\[1/\sec=\cos,~~~so,\]\[tanx+secx=(cosx) \times tanx\]\[tanx+secx=cosx~tanx\]\[tanx=sinx/cosx~~~AND~~~secx=1/cosx~~~~So,\]\[\frac{Sinx}{Cosx}+\frac{1}{Cosx}=cosx \times \frac{Sinx}{Cosx}\]\[\frac{Sinx+1}{Cosx}=Sinx\]\[Sinx+1=SinxCosx\]\[\frac{Sinx}{Sinx}+\frac{1}{Sinx}=\frac{SinxCosx}{Sinx}\]\[1+Cscx=Cosx\]\[1=Cosx-Cscx\]

OpenStudy (anonymous):

I think this is an open equation, it's not true or false, you would want to solve for x.

OpenStudy (anonymous):

no, I mean how can I prove that the right side of the equation is equal to the left side using the fundamental identities

OpenStudy (anonymous):

please help me.

OpenStudy (amistre64):

change them to sin cos identities to see how to manipulate it easier

OpenStudy (anonymous):

I already did that but still I'm confused.

OpenStudy (anonymous):

please, if anyone knows this identity please hep me

OpenStudy (amistre64):

lets make sure it is an identity first :) http://www.wolframalpha.com/input/?i=tan%28x%29%2Bsec%28x%29%3D1%2F%28sec%28x%29tan%28x%29%29 so the issue is, that it is NOT an identity to start with; so no matter how much you try to manipulate it it will NEVER work out as an identity

OpenStudy (amistre64):

we can solve for some specific values of x that will work out, but an identity has to work for ALL values of x.

OpenStudy (anonymous):

so it means that the equation given was not qualified to be proved as equal identities?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

youre welcome